Normal Distributions:

Finding Values A particular fruit's weights are normally distributed, with a mean of 381 grams and a standard deviation of 13 grams. The heaviest 3% of fruits weigh more than how many grams?

Give your answer to the nearest gram.
Answer #1

Given that,

mean = = 381

standard deviation = =13

Using standard normal table,

P(Z > z) = 3%

= 1 - P(Z < z) = 0.03

= P(Z < z ) = 1 - 0.03

= P(Z < z ) = 0.97

z = 1.88 (using standard normal (Z) table )

Using z-score formula  

x = z * +

x= 1.88 *13+381

x= 405.44=405